OpenStudy (anonymous):

1. In your own words, explain how to find the first quartile of a set of data. 2. What is the interquartile range of the following speeds recorded by the state police? (53, 55, 55, 55, 57, 57, 60, 60, 62, 63, 63, 64, 67, 68, 68, 70, 72, 73) 3. What is the median of the first half of a data set called? 4. What is the median of the second half of a data set called? 5. Explain how to find the IQR?

OpenStudy (amistre64):

so how would you explain how to find the 1st Q?

OpenStudy (amistre64):



OpenStudy (amistre64):

ideally, a quartile is a position within the data set; so we divide it in quarters and find the 1st quarter

OpenStudy (campbell_st):

quartiles divide a data set into 4 equal parts steps 1. find the median. The position is found by using (n + 1)/2 n = number of data scores in your example there are 18 scores Position (18 +1)/2 = 9.5 the median is between th 9th and 10th score (62 + 63)/2 = 62.5 the median is 62.5 Lower Quartile of Q1. find the middle of the lower half of the data 53, 55, 55, 55, 57, 57, 60, 60, 62 9 scores the median position is (9 + 1)/2 = 5 the 5th score is the lower quartile 57 Upper Quartile of Q3 63, 63, 64, 67, 68, 68, 70, 72, 73 9 scores so Q3 is in position 5 Q3 = 68 The interquartile range is the distance between Q3 and Q1 IQR = 68 - 57 IQR = 11